OpenStudy (anonymous):

is the sin120 degrees -1/2?

OpenStudy (anonymous):

\[\sin 120=\sin (180-60)=\sin 60=\sqrt {3}\]

OpenStudy (anonymous):

/2

OpenStudy (anonymous):

just remember that on the unit circle, sin is your y-coordinate cos is your x-coordinate.

OpenStudy (zehanz):

Still on the top-half of the unit circle, so sin is positive there...

OpenStudy (anonymous):

is this a complete question

OpenStudy (anonymous):

what is it looking for

OpenStudy (anonymous):

it's basically a yes/no question...

OpenStudy (anonymous):

it just said find each trigonometric ratio and gave me sin120.

OpenStudy (anonymous):

oh i see, so since sin is the y axis, it would be sqrrt3/2

OpenStudy (anonymous):

correct... sin is the y-coordinate, not the y axis...